### v1

# Controllable diatomic molecular quantum thermodynamic machines

Abstract
We present quantum heat machines using a diatomic molecule modelled by a $q$-deformed potential as a working medium. We analyze the effect of the deformation parameter and other potential parameters on the work output and efficiency of the quantum Otto and quantum Carnot heat cycles. Furthermore, we derive the analytical expressions of work and efficiency as a function of these parameters. Interestingly, our system reveals different operational phases for the range of parameter regimes considered, such as a quantum heat engine. In addition, the efficiency of the quantum Otto heat engine is seen to be tunable by the deformation parameter. Our findings provide useful insight for understanding the impact of anharmonicity on the design of quantum thermal machines.
I Introduction
Thermal machines have been at the heart of thermodynamics since the early days of the field [1]. They are often harnessed to provide mechanical energy from thermal energy. Macroscopic motors, refrigerators, and heat pumps are notable examples of classical machines whose dynamics follow the laws of classical physics [1]. The advances in nanofabrication have led to the development of smaller devices. On this scale, the quantum effects dominate, and the laws must be considered. Since the 1950s, there have been intense theoretical endeavours in comprehending the interplay of thermodynamics and quantum physics following the seminal paper Scovil and Schulz-DuBois on maser heat engines [2].
Quantum heat engines produce work using quantum matter as a working substance and exhibit exotic properties. They have been shown to perform better compared to their classical counterpart [3]. Several quantum mechanical models have been proposed as a working medium in recent times, for example, the harmonic oscillator potential [4], infinite square well potential [4], spin systems [5, 6, 7, 8], stanene [9], strained-graphene [10], Dirac particles [11, 12], two-level system [13], multi-level system [14], a continuum working medium [15], photon gas [16], diatomic molecules [17], Posch-Teller potential [18, 19], among others. In addition, the past decade has witnessed several experimental implementations of quantum thermal machines [20, 3, 21, 22, 23, 24, 25].
Moreover, quantum systems that exhibit anharmonicity, such as molecular systems [26, 27], superconducting qubits [28], are of great interest in the development of quantum technologies. Anharmonic effects have been shown to be useful in several settings [29], but most quantum systems are often treated with approximate harmonic theoretical descriptions, for instance in the case of a single molecule [30, 31, 32, 33, 34, 35, 36, 37]. The harmonic approximation is valid when atomic displacements are rather limited [38], and recent experiments are beginning to probe the limits of this model in a plethora of processes and contexts [39, 40, 41]. However, most quantum thermal machines assume that their working mediums are often harmonic in nature [42, 20, 3], but recent experiments show the possibility of trapping molecules in anharmonic traps [29]. Therefore, it becomes crucial to develop quantum thermal machines with genuine anharmonicity to see how they impact the performance of thermal machines. In this regard, a Carnot-like engine with a diatomic molecule has recently been proposed by Oladimeji et al. [17]. Here, we extend this to carry out a full quantum thermodynamic description of a quantum thermal machine whose working medium is a tunable Morse potential, which gives us the “degree of freedom” to tune the behaviour of the model. The Morse potential is a well-known diatomic molecular potential that captures the effect of anharmonicity and real bond-breaking [26]. We investigate the anharmonic effects on the performance of a quantum thermal machine in an analytically feasible model using a generalized ($q$-deformed) Morse potential as the working substance. The choice of this model is inspired by the fact the presence of the $q$-parameter gives us tunable control in our setup. This $q$-deformed physical system is typically described by q-deformed potentials which is reduced to their normal equivalent when $q\rightarrow 1$. The q-deformed hyperbolic potentials first appeared in the seminal works of Arai [43, 44]. This class of model potentials has since found applications in several areas of research, notably the study of the phonon spectrum in ${}^{4}\text{He}$ [45], multi-atomic molecules [46], oscillatory-rotational spectra of diatomic [47], the description of the electronic conductance in disordered metals and doped semiconductors [48] amongst many other[49, 50, 51]. A recent study showed that statistical complexity on the information-theoretic measure of diatomic molecules is strongly influenced by the potential deformation parameter $q$ [52]. Therefore, it is the goal of this study to extend the application of such a physical system to quantum thermodynamics.
In this paper, we study the performance of a quantum thermal machine using a diatomic molecule as a working substance. We show that the presence of the deformation parameter and other potential parameters can be used to regulate the behaviour of the system such that different operational phases are revealed in different parameter regimes. This article is organized as follows. We introduce the diatomic molecule model in Sec. II. In Sec. III, we present the theory and analysis of the quantum thermodynamic Carnot and the quantum Otto cycle of the controllable molecule model. In Sec. IV, we present concluding remarks and discuss experimental feasibility
II The diatomic molecule
We begin by describing the model and previous results from the literature that are central to our analysis.
Figure 1: Plot of the q-deformed Morse potential for $q\!\in\!\{0.4r_{e},~{}0.5r_{e},~{}r_{e}\}$, where $q\rightarrow 1$ corresponds to the standard Morse potential. Parameters used: $\mathcal{D}_{\text{e}}\!=\!10r_{e}$, and $\alpha\!=\!2r_{e}$.
The Hamiltonian describing the diatomic molecule system reads [53, 50, 52, 54, 55],
$H$ $=$ $\frac{p_{x}^{2}}{2\mu r_{e}^{2}}+\mathcal{V}_{q}(x),$ (1)
$=$ $\frac{p_{x}^{2}}{2\mu r_{e}^{2}}+\mathcal{D}_{e}(e^{-2\xi x}-2qe^{-\xi x}),$
where the first term is the kinetic energy operator with $p_{x}\!=\!-i\hbar d/dx$ as the momentum operator, and $d/dx$ is the differential operator. The second term is the deformed Morse potential model $\mathcal{V}_{q}(x)$, where $x\!=\!(r-r_{e})/r_{e}$, and $\xi=\alpha\,r_{e}.$ The first term of the potential corresponds to the repulsive part, and the second terms correspond to the attractive part of the potential. Here, $\mathcal{D}_{e}>0$ is the potential strength or dissociation energy, it quantifies the depth of the potential well at the equilibrium distance, $r_{e}$ is the equilibrium bond length, $\alpha$ determines the width of the potential well, and $q$ is a dimensionless parameter for controlling the shape of the potential (see, Fig. (1). The deformation parameter $q$, it is a real number $q\in[0,1]$. We recover the standard Morse potential [26] when $q=1$. This Morse potential has been noted to be a suitable potential model for modeling the inter-atomic interaction in a diatomic molecule. This is because it accounts for the impact of bond breaking, such as the existence of unbound states. More so, it also accounts for the anharmonicity of real bonds. On expansion of $\mathcal{V}_{q}(x)$ from Eq. (1) in a power series of $x$, we get
$\mathcal{V}_{q}(x)\approx D_{e}(1-2q)+2\xi D_{e}(q-1)x+\xi^{2}D_{e}\left(2-q\right)x^{2}+O\left(x^{3}\right)$ (2)
It is interesting to note that in contrast to the standard Morse potential, $q$-deformed Morse potential cannot be approximated by the quantum harmonic oscillator except in the limit of $q\rightarrow 1$, $\mathcal{V}_{q}(x)\!\approx\!-D_{e}+\xi^{2}D_{e}x^{2}+O\left(x^{3}\right)$ in the neighbourhood of its minimum. Hence, we can see that for small values of $x$ the Morse potential, Eq. (1), can be approximated by a harmonic oscillator with spring constant $k=2\xi^{2}D_{e}$, and $V_{q}(x)\!\approx\!-D_{e}+k/2x^{2}$, this was illustrated by Leonard and Deffner [56]. It is observed that the Morse potential significantly deviates from the harmonic approximation as the distance from $x\!=\!0$ increases. This allows us to model bond weakening and dissociation with diverging nuclear separation. Moreover, the presence of the deformation reveals other features such as symmetry distortion [52]. This is particularly evident in the regime where the $q$ parameter is very small. The consequence of these properties introduces interesting features into the quantum heat engine performance, as we will discuss shortly in Sec. III.
The solutions of the Schrodinger equation for the Hamiltonian, Eq. (1), yield the energy eigenvalue (see Appendix A for details):
$\mathcal{E}_{n}=-\xi^{2}p\left(-n+\lambda q-\frac{1}{2}\right)^{2},$ (3)
where $n\!=\!0,1,2,3,...,n_{\text{max}}$, $\lambda\!=\!\sqrt{\frac{D_{e}}{\xi^{2}p}}$, and $p\!=\!\frac{\hbar^{2}}{2\mu r_{e}^{2}}$. The energy level associated with the $q$-deformed Morse oscillator is bounded from above by $q$ with the maximum quantum number given as $n_{\text{max}}\leq(\lambda q-1/2)$. The equilibrium thermodynamic behaviour of the $q$-deformed Morse potential can be determined from the canonical partition function,
$\mathcal{Z}=\text{Tr}[\text{exp}(-\beta H)]=\sum\limits_{n=0}^{n=\infty}\text{exp}(-\beta\mathcal{E}_{n}),$ (4)
where $\beta\!=\!1/(k_{\mathrm{B}}T)$ is the inverse temperature $T$ with the Boltzmann constant $k_{\mathrm{B}}$. In the energy representation, it is straightforward to calculate the partition function using the $q$-deformed Morse potential eigen energies (3), we obtain
$\mathcal{Z}=\frac{\sqrt{\pi}\text{erfc}\left(\frac{1}{2}(1-2\lambda q)\sqrt{\beta\xi^{2}(-p)}\right)}{2\sqrt{\beta\xi^{2}(-p)}}.$ (5)
The thermal state density matrix can be determined from the eigenstate and the partition function as,
$\rho=\sum\limits_{n}\mathcal{P}_{n}|n\rangle\langle n|=\sum\limits_{n}(\text{exp}(-\beta\mathcal{E}_{n})/\mathcal{Z})|n\rangle\langle n|,$ (6)
where $\mathcal{P}_{n}$ is the occupation probability of the $n^{\text{th}}$ eigenstate and $\mathcal{E}_{n}$ is the $n^{\text{th}}$ eigen energy of the working medium, $q$-deformed Morse potential.
Hence, the total internal energy reads $\mathcal{U}\!=\!\sum\limits_{n}\mathcal{P}_{n}\mathcal{E}_{n}$, and the infinitesimal change yields $d\mathcal{U}\!=\!\sum\limits_{n}(\mathcal{E}_{n}d\mathcal{P}_{n}+\mathcal{P}_{n}d\mathcal{E}_{n})$. Thus, the first law of thermodynamics can be deduced as,
$d\mathcal{U}\!=\!d\mathcal{Q}+d\mathcal{W},$ (7)
where $d\mathcal{Q}\!=\!\sum\limits_{n}\mathcal{E}_{n}d\mathcal{P}_{n}$ and $d\mathcal{W}\!=\!\sum\limits_{n}\mathcal{P}_{n}d\mathcal{E}_{n}$. The internal energy, $\mathcal{U}$, is the state function, whereas heat $\mathcal{Q}$ and work $\mathcal{W}$ are path-dependent functions. We remark that recent studies have focused on analyzing the equilibrium thermodynamic properties of the $q$-deformed Morse potential within the framework of superstatistics [57, 54] and the Euler-Maclaurin method [55].
III Diatomic molecular quantum heat engine
An aggregate of different quantum thermodynamic processes and the number of strokes per cycle yields several kinds of quantum heat engines [58, 59]. In this section, we present the thermodynamics analysis of the performance of the diatomic molecule engine for two distinct thermodynamic cycles, namely the Carnot and Otto cycles.
III.1 Quantum Carnot Cycle
[…]
$\mathcal{W}^{\text{QC}}=Q_{\text{h}}+Q_{\text{c}}=(T_{\text{h}}-T_{\text{c}})\left[S(\text{B})-\mathcal{S}(\text{D}))\right],$ (8)
where $\mathcal{S}(i)$ is the entropy given as
$\mathcal{S}(i)=-k\sum\frac{\exp[-\beta_{i}\mathcal{E}_{n}(i)]}{\mathcal{Z}(i)}[-\beta_{i}\mathcal{E}_{n}(i)-\text{ln}\mathcal{Z}(i)],$ (9)
[…]
$\mathcal{S}_{h(c)}=k\log\left(\frac{\sqrt{\pi}\text{Erfc}\left(\gamma_{1}^{h(c)}\right)}{2\sqrt{\beta_{h(c)}\xi_{h(c)}^{2}(-p)}}\right)-\frac{1}{2}k\left(\Gamma_{1}-1\right),$ (10)
where $\gamma_{1}^{h(c)}\!=\!\frac{1}{2}(1-2\lambda_{h(c)}q)\sqrt{\beta_{h(c)}\xi_{h(c)}^{2}(-p)},\gamma_{2}^{h(c)}=(2\lambda_{h(c)}q-1)\sqrt{\beta_{h(c)}\xi_{h(c)}^{2}(-p)}$, $\gamma_{3}^{h(c)}\!=\!\frac{1}{4}\beta_{h(c)}\xi_{h(c)}^{2}p(1-2\lambda_{h(c)}q)^{2}$, and $\Gamma_{1}=\left(\gamma_{2}^{h(c)}e^{\gamma_{3}^{h(c)}}\right)/\left(\sqrt{\pi}\text{Erfc}\left(\gamma_{1}^{h(c)}\right)\right)$.
Figure 2: A pictorial representation of a Carnot thermal machine using a $q$-deformed anharmonic oscillator as a working substance. The engine cycle consists of two adiabatic strokes ($B\to C$ and $D\to A$) and two isothermal strokes, $A\to B$ at $T_{\text{h}}$ and $C\to D$ at $T_{\text{c}}$, with $T_{\text{h}}>T_{\text{c}}$.
Figure 3: Quantum Carnot engine: (a) The heat exchanges of the working substance with the hot reservoir, $\mathcal{Q}_{\text{h}}$ vs $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$ (b) The heat exchanges of the working substance with the cold reservoir, $\mathcal{Q}_{\text{c}}$ vs $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. (c) The Work output $\mathcal{W}$ vs $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. Parameters used: $\alpha_{h}=2.236r_{e}$, $\alpha_{c}=r_{e}$, $T_{c}\!=\!2r_{e}$ and $T_{h}=10r_{e}$
The reversibility condition is defined as [15]:
$\mathcal{E}_{\text{n}}(\text{C})-\mathcal{E}_{\text{m}}(\text{C})=\frac{T_{\text{c}}}{T_{\text{h}}}[\mathcal{E}_{\text{n}}(\text{B})-\mathcal{E}_{\text{m}}(\text{B})],$ (11)
where $\mathcal{E}_{n(m)}$ is the $q$-deformed Morse potential energy eigenvalue. The reversibility condition is satisfied independent of the states $n(m)$ and the variation in the width ($\alpha$) of the $q$-deformed Morse potential while other parameters such as the deformation parameter $q$, the dissociation energy $\mathcal{D}_{e}$, and the equilibrium bond length $r_{e}$ are kept constant during the cycle. Thus, if $\sqrt{\mathcal{D}_{e}^{\text{C}}/\alpha_{\text{C}}^{2}}\!=\!\sqrt{\mathcal{D}_{e}^{\text{B}}/\alpha_{\text{B}}^{2}}$, we obtain
$\frac{\alpha_{\text{C}}^{2}}{\alpha_{\text{B}}^{2}}=\frac{T_{c}}{T_{h}},$ (12)
as the reversibility condition. Equation (12) shows that the reversibility condition is satisfied by taking appropriate values of the width of the $q$-deformed Morse potential. Now, using the entropies within the reversibility condition, the work done for a quantum Carnot engine whose working medium is a $q$-deformed Morse oscillator can be expressed as;
$\mathcal{W}^{\text{QC}}=\frac{k(T_{\text{c}}-T_{\text{h}})\left(\text{Erfc}\left(\gamma_{1}^{c}\right)\left(2\sqrt{\pi}\text{Erfc}\left(\gamma_{1}^{h}\right)\left(\log\left(\frac{\text{Erfc}\left(\gamma_{1}^{c}\right)}{\sqrt{\beta_{c}\xi_{c}^{2}(-p)}}\right)-\log\left(\frac{\text{Erfc}\left(\gamma_{1}^{h}\right)}{\sqrt{\beta_{h}\xi_{h}^{2}(-p)}}\right)\right)+\gamma^{h}_{2}e^{\gamma_{3}^{h}}\right)-\gamma_{2}^{c}e^{\gamma_{3}^{c}}\text{Erfc}\left(\gamma_{1}^{h}\right)\right)}{2\sqrt{\pi}\text{Erfc}\left(\gamma_{1}^{c}\right)\text{Erfc}\left(\gamma_{1}^{h}\right)}.$ (13)
The corresponding efficiency reads,
$\eta^{\text{QC}}=\frac{W^{\text{QC}}}{\mathcal{Q}_{\text{h}}}=\frac{(T_{\text{h}}-T_{\text{c}})(\mathcal{S}_{\text{h}}-\mathcal{S}_{\text{c})}}{T_{\text{h}}(\mathcal{S}_{\text{h}}-\mathcal{S}_{\text{c})}}=1-\frac{T_{\text{c}}}{T_{\text{h}}}.$ (14)
In Fig. (3), we present the numerical results of the heat exchanges and the total work extracted against the control/deformation parameter ($q/r_{e}$) and the dissociation energy ($D_{e}/r_{e}$). Fig. 3 (a) shows that the input heat is positive, that is $\mathcal{Q}_{\text{h}}>0$ while Fig. 3(b) demonstrate that the output heat is always negative $\mathcal{Q}_{\text{c}}<0$ within the parameters range. In Fig. 3(c), we observe an increase in work output as both $q/r_{e}$ and dissociation energy $D_{e}/r_{e}$ increase. Both parameters enhance the work output but have a stronger effect when they are at their maximum. This demonstrates that the width of the well and the hot bath temperature play a crucial role in enhancing the work output. Thus, considering that the total work output is positive, we conclude that the system operates as a quantum heat engine in this regime. The efficiency of the quantum Carnot $q$-deformed Morse oscillator heat engine is 0.8 at temperatures that satisfy the reversibility condition.
III.2 Quantum Otto Cycle
This consists of four strokes; two quantum isochoric and two quantum adiabatic processes (see Fig.4). In the quantum adiabatic process, $B\to C$ and $D\to A$, the occupational probability of the eigenstate is invariant, that is, $\mathcal{P}_{n}(A)\!=\!\mathcal{P}_{n}(D)$ and $\mathcal{P}_{n}(B)\!=\!\mathcal{P}_{n}(C)$. This is because it satisfies the general adiabatic condition $d\mathcal{P}_{n}\!=\!0$. During this process, the system is isolated and does not exchange heat with the environment while the work is non-zero.
Figure 4: A pictorial illustration of an Otto machine The engine cycle consists of two adiabatic strokes ($B\to C$ and $D\to A$) where it is decoupled from the thermal baths and two isochoric strokes ($A\to B$ and $C\to D$) where the engine is coupled to two thermal baths at temperatures $T_{\text{h}}$ and $T_{\text{c}}$, with $T_{\text{h}}>T_{\text{c}}$.
The quantum isochoric process represents $A\to B$ and $C\to D$. In this process, while heat is exchanged between the system and the reservoir, no work is done. Hence, the energy of the eigenvalue remains constant, that is, $\mathcal{E}^{\text{h}}\!=\!\mathcal{E}(B)\!=\!\mathcal{E}(A)$ and $\mathcal{E}^{\text{c}}\!=\!\mathcal{E}(D)\!=\!\mathcal{E}(C)$. Furthermore, in the isochoric process, $\mathcal{P}_{n}$ and $\mathcal{S}$ change until equilibrium is reached. Heat is released ($\mathcal{Q}_{\text{c}}<0$) when the entropy decreases in stroke $C\to D$, while heat is absorbed ($\mathcal{Q}_{\text{h}}<0$) when the entropy increases for $A\to B$, satisfying the Boltzmann distribution [4]. The heat absorbed by the working medium during the heating process is
[…]
where $\mathcal{E}^{\text{h}}$ is the $n^{th}$ energy of the system from $B\to C$. Furthermore, the heat released during the cooling process is
[…]
where $\mathcal{E}^{\text{c}}$ is the $n^{\text{th}}$ energy of the cooling process of the system from $D\to C$. The work done for the Otto heat engine is the sum of $\mathcal{Q}_{\text{h}}$ and $\mathcal{Q}_{\text{c}}$ given as [4, 60];
$\mathcal{W}^{\text{QO}}=\sum\limits_{n}\left[\mathcal{E}^{h}_{n}-\mathcal{E}^{\text{c}}_{n}\right]\left[\mathcal{P}_{n}(B)-\mathcal{P}_{n}(D)\right],$ (17)
and the corresponding efficiency of the $q$-deformed Morse oscillator quantum heat engine is [58, 13]
$\eta_{otto}=\frac{\mathcal{W}^{\text{QO}}}{\mathcal{Q}_{\text{h}}}.$ (18)
In what follows, we present the performance of three different scenarios for realizing a quantum Otto engine cycle using a diatomic molecule as the working medium.
III.2.1 Changing width of the potential well
Now consider the case in which the width of the potential well $\alpha/r_{e}$ and the temperature $T/r_{r_{e}}$ are varied during the cycle. Using the energy eigenvalue expression and Eq. (16), we obtain the analytical expression for the input heat of the quantum Otto engine as,
$\mathcal{Q}^{\text{QO}}_{\text{h}}$ $=$ $\frac{1}{2\beta_{\text{h}}\sqrt{\pi}}\left(\frac{2\gamma_{1}^{\text{h}}e^{\Lambda_{1}^{\text{h}}}}{\text{erfc}\left(\gamma_{1}^{\text{h}}\right)}+\sqrt{\pi}\left(-\frac{\beta_{\text{h}}\xi_{\text{h}}^{2}}{\beta_{\text{c}}\xi_{\text{c}}^{2}}+\Lambda_{3}^{\text{ch}}\right)\right)$ (19)
$+$ $\frac{1}{2\sqrt{\pi}}\frac{\Lambda_{2}^{\text{c}}e^{\Lambda_{1}^{\text{c}}}}{\sqrt{\beta_{\text{c}}\xi_{\text{c}}^{2}(-p)}\text{erfc}\left(\gamma_{1}^{\text{c}}\right)},$
where $\Lambda_{1}^{\text{c}},\Lambda_{1}^{\text{h}},\Lambda_{2}^{\text{ch}},\Lambda_{3}^{\text{ch}},\text{and}\,2\gamma_{1}^{\text{h}}$ are dimensionless parameters defined in Eq. (35) of Appendix (B). Similarly, the expression for the output heat $\mathcal{Q}^{\text{QO}}_{\text{c}}$ reads
$\mathcal{Q}^{\text{QO}}_{\text{c}}$ $=$ $\frac{1}{2\sqrt{\pi}}\left(\frac{2\gamma_{1}^{\text{c}}e^{\Lambda_{1}^{\text{c}}}}{\beta_{\text{c}}\text{erfc}\left(\gamma_{1}^{\text{c}}\right)}+\frac{\Lambda_{2}^{\text{hc*}}e^{\Lambda_{1}^{\text{h}}}}{\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}\text{erfc}\left(\gamma_{1}^{\text{h}}\right)}\right)$ (20)
$+$ $\frac{1}{2}\left(\frac{1}{\beta_{\text{c}}}+\Lambda_{4}^{ch}\right),$
where the dimensionless parameters $\Lambda_{2}^{\text{hc*}}2\gamma_{1}^{\text{c}},\gamma_{1}^{\text{h}},\text{and}\Lambda_{4}^{ch}$ are defined explicitly in Eq. (35), and (43) of Appendix (B). Using Eq. (17), we obtain the work output as follows:
$\mathcal{W}^{\text{QO}}=\frac{1}{2\sqrt{\pi}}\left(\frac{\Lambda_{10}+\Lambda_{11}\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}}{\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}\text{{\hbox{\beta}}c}\beta_{\text{h}}\xi_{\text{h}}^{2}\text{erfc}\left(\gamma_{1}^{\text{h}}\right)}-\frac{p\mathcal{R}_{0}}{\mathcal{R}_{1}}\right)$ (21)
where the dimensionless parameters $\Lambda_{5}^{\text{ch}}$, $\Lambda_{6}^{\text{ch}}$, $\Lambda_{7}^{\text{ch}}$, $\Lambda_{8}^{\text{ch}}$, $\Lambda_{9}^{\text{h}}$, $\Lambda_{9}^{\text{c}}$, $\Lambda_{10},\Lambda_{11},\Lambda_{12}$, $\mathcal{R}_{0}\!=\!\left(\Lambda_{12}+\Lambda_{13}\right)$, and $\mathcal{R}_{1}\!=\!\left(\beta_{\text{c}}\xi_{\text{c}}^{2}(-p)\right)^{3/2}\text{erfc}\left(\gamma_{1}^{\text{c}}\right)$ are defined explicitly in Eqs. (35), (43) and (48) of Appendix (B). Finally, using Eq. (18) and (47), an explicit analytical expression for efficiency is obtained (see Eq. (59) in Appendix (B)).
Figure 5: Quantum Otto engine with changing width of the potential well: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$, (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ (c) Work output, $\mathcal{W}$, and (d) Efficiency, $\eta$ as a function of $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. Parameters used: $\alpha_{h}\!=\!2.236\,r_{e}$, $\alpha_{c}\!=\!r_{e}$, $T_{c}\!=\!2\,r_{e}$ and $T_{h}\!=\!10\,r_{e}$.
We present the Otto engine thermodynamic figure of merit quantities in Fig. (5). Figures 5(a) and (b) show that $\mathcal{Q}_{h}\!>\!0$ and $\mathcal{Q}_{c}\!<\!0$ are satisfied within a wide range of well width and potential deformation. In Fig. (5) (c), we observe an increase in the work performed as both $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$ increase. This is particularly seen especially moving from left (low) $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$ towards the top-right region. The numerical result generally shows a positive value for the output work of the Otto heat engine. The value is very high in the region $0.95r_{e}<q<0.99r_{e}$ and $13.9\,r_{e}\!<\!\mathcal{D}_{e}\!<\!15\,r_{e}$. Although a slight flip to a lower value is observed around $q\!=\!r_{e}$. Based on the magnitudes of $\mathcal{Q}_{\text{h}}\!>\!0$ and $\mathcal{Q}_{\text{c}}\!<\!0$, we conclude that in this parameter regime considered, the system can operate as a quantum heat engine. In Fig. 5 (d), we present the efficiency of the quantum Otto heat engines as a function of $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. We observe that for the parameter range considered, the efficiency of the Quantum Otto engine is lower than the efficiency of the Quantum Carnot Heat engine ($\eta_{C}\!=\!0.8$), which is in agreement with the literature [62, 60]. Focusing on Otto engine efficiency, we see that as $q\rightarrow 1$, the efficiency decreases; this is in agreement with earlier studies by Ozaydin et al. [63] that focus on the $q$-deformed quantum oscillator as a working substance. Furthermore, this demonstrates that the deformation introduced is able to tune the efficiency and, by extension, other important figures of merit of our system.
Figure 6: Quantum Otto cycle with the changing deformation parameter: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$ (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ (c) Total work output, $\mathcal{W}$ (d) Efficiency $\eta$ as function of $\mathcal{D}_{e}/r_{e}$ and $\alpha/r_{e}$. Parameters used: $q_{c}=0.8r_{e}$, $q_{h}=r_{e}$, $T_{h}=10r_{e}$, and $T_{c}=2r_{e}$.
III.2.2 Changing the strength of the deformation parameter
Here, the strength of the deformation parameter $(q/r_{e})$ and the temperature $T/r_{e}$ vary during the cycle while the dissociation energy ($\mathcal{D}_{e}/r_{e}$) and the width of the potential well ($\alpha$) are kept constant. The work output, using Eq.(17), is
$\mathcal{W}^{\text{QO}}=\sum\limits_{n}\left[\mathcal{E}^{\text{h}}_{n}(q_{\text{h}})-\mathcal{E}^{\text{c}}_{n}(q_{\text{c}})\right]\left[\mathcal{P}_{n}(B)-\mathcal{P}_{n}(\text{D})\right],$ (22)
where $\mathcal{E}^{\text{h}}_{n}(q_{\text{h}})$, and $\mathcal{E}^{c}_{n}(q_{\text{c}})$ are the $n^{\text{th}}$ energy levels associated with the two isochoric processes which imply constant energy levels. Also, for the quantum adiabatic strokes during which the occupation probabilities of energy eigenstates remain unchanged. When the strength of the deformation parameter is varied, it is important to note that the Carnot reversibility condition is not satisfied [4].
In Fig. (6), we show the figure of merit of the performance of the quantum Otto engine using a diatomic molecule as its working substance for the case of varying deformation strength. Figures (6) (a) and (b) show the heat exchanges between the working substance and the heat reservoirs during the hot and cold isochoric process, respectively. Analyzing the magnitude of $\mathcal{Q}_{\text{h}}>0$ and $\mathcal{Q}_{\text{c}}<0$, as shown in Figs. (6) (a) and (b), we conclude that in this parameter regime considered, the system operates as a quantum heat engine. In Fig. (6) (c) show that the work output is high in the region of low $\alpha$ and high $\mathcal{D}_{e}/r_{e}$. This is particularly seen as we move from right (low) to left (high) $\mathcal{D}_{e}/r_{e}$ and $\alpha/r_{e}$ towards the top-region region. The plot generally shows a positive value of the work output for a wide range of parameters. This is also very intense in the region $r_{e}\!<\alpha<\!2r_{e}$ and $13\,r_{e}<\mathcal{D}_{e}<15\,r_{e}$. In Fig. (6) (d), the efficiency is seen to be predominantly high in the region of low $\alpha/r_{e}$ and high $\mathcal{D}_{e}$.
III.2.3 Changing the strength of the dissociation energy
We consider the case of constructing a quantum Otto engine by varying the dissociation energy $(\mathcal{D}_{e}/r_{e})$ and the temperature $T/r_{e}$. The work performed in this case using (17) is
$\mathcal{W}^{\text{QO}}=\sum\limits_{n}\left[\mathcal{E}^{h}_{n}(\mathcal{D}_{e}^{\text{h}})-\mathcal{E}^{\text{c}}_{n}(\mathcal{D}_{e}^{\text{c}})\right]\left[\mathcal{P}_{n}(B)-\mathcal{P}_{n}(D)\right].$ (23)
Analyzing Figs. (7) (a) and (b) reveal that $\mathcal{Q}_{\text{h}}>0$ and $\mathcal{Q}_{\text{c}}<0$, over the range of parameters. The positive value of the total work output, Fig. (7) (c), implies that the system functions as a quantum heat engine. The work output is low in the high regime $\alpha/r_{e}$ and increases as $q/r_{e}$ increases. In Fig. (7) (d), the efficiency is observed to be predominantly high (low) in the region of high (low) $q/r_{e}$($\alpha$) and vice versa.
Figure 7: Quantum Otto cycle for varying dissociation energy strength: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$ vs $\alpha/r_{e}$ and $q/r_{e}$. (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ vs $\alpha/r_{e}$ and $q/r_{e}$. (c) Work output of Otto engine $\mathcal{W}$ vs $\alpha/r_{e}$ and $q/r_{e}$. (d) Efficiency $\eta$ vs $\alpha/r_{e}$ and $q/r_{e}$. Parameters used:$\mathcal{D}_{e}^{\text{c}}=5r_{e}$, $\mathcal{D}_{e}^{\text{h}}=10r_{e}$, $T_{h}=10r_{e}$, and $T_{c}=2r_{e}$.
IV Conclusions
We presented a theoretical proposal for a quantum thermal machine with the $q$-deformed Morse potential as a working medium. We discussed in detail the different regimes of operation for our heat engine. These modes of operation can be modulated by appropriate adjustment of the anharmonic potential parameters. Furthermore, we present analytical expressions for the work done and the efficiency of the heat engine. We find that these figures of merits depend on the variations of the potential parameters. It is also noted that the performance of both the engine and the refrigerator can be regulated or controlled by adjusting the deformation parameter and other system parameters. The potential platform for the realization of our model is a laser-cooled trapped ion as a microscopic heat machine, proposed by Gelbwaser-Klimovsky et al. [64]. However, in our context, the trap will be anharmonic, which will be more suitable for trapping a diatomic molecule. The experimental considerations for the realization of a tunable anharmonic trap have been presented by Home et al. [29]. Our findings may inspire several developments in the emerging field of quantum technologies such as the design of quantum thermal machines based on anharmonic models.
Acknowledgements
COE and EPI gratefully acknowledge the support provided by the Tertiary Education Trust Fund-Institutional Based Research (TETFUND-IBR) grant scheme with grant number: NOUN/DRA/TETFUNDAW/VOLI. COE and NA acknowledge support from the Long Term Research Grant Scheme (LRGS) with grant number LRGS/1/2020/UM/01/5/2 (9012-00009), provided by the Ministry of Higher Education of Malaysia (MOHE).
Appendix A Solutions of the Schrödinger Equation for the q-deformed Morse Potential
The Schrodinger equation in one-dimensional has the form
$\left(-\frac{\hbar^{2}}{2\mu r_{e}^{2}}\frac{d^{2}}{dx^{2}}+\mathcal{V}_{q}(x)\right)\psi_{n}(x)=\mathcal{E}_{n}\psi_{n}(x)$ (24)
Considering the q-deformed Morse potential as
$\mathcal{V}_{q}(x)=\mathcal{D}_{e}(e^{-2\xi x}-2qe^{-\xi x})$ (25)
Substituting Eq. (25) in Eq. (24), we obtain
$\left(\frac{d^{2}}{dx^{2}}-\frac{2\mu\mathcal{D}_{e}}{\hbar^{2}}e^{-2\xi x}+\frac{4\mu q\mathcal{D}_{e}}{\hbar^{2}}e^{-\xi x}+\frac{2\mu}{\hbar^{2}}\mathcal{E}_{n}\right)\psi_{n}(x)=0$ (26)
Assuming $y=\delta e^{-\xi x}$ , Eq. (26) takes the form
$\left(\frac{d^{2}}{dy^{2}}+\frac{1}{y}\frac{d}{dy}+\frac{-\text{s}^{2}}{y^{2}}+\frac{\delta q}{2y}-\frac{1}{4}\right)\psi_{n}(y)=0$ (27)
where $s=\sqrt{-\frac{2\mu r_{e}^{2}\mathcal{E}_{n}}{\hbar^{2}\xi^{2}}}$, $\delta=\sqrt{\frac{8\mu r_{e}^{2}\mathcal{D}_{e}}{\hbar^{2}\xi^{2}}}$ Looking at the behaviour of the wave function at the origin and at infinity, we can take the following ansatz as
$\psi(y)=\text{e}^{-\frac{y}{2}}y^{s}\mathcal{F}(y)$ (28)
Substituting the wave function ansatz (28) into Eq. (27), this yields
$y\frac{d^{2}}{dy^{2}}\mathcal{F}(y)+(2s+1-y)\frac{d}{dy}\mathcal{F}(y)-\left(s+\frac{1-\delta q}{2}\right)\mathcal{F}(y)=0,$ (29)
whose solutions are the confluent hypergeometric functions as
$\mathcal{F}(y)=\mathcal{N}\mathcal{F}\left(s+\frac{1-\delta q}{2},2s+1;y\right)$ (30)
Thus, we obtain the quantum condition
$s+\frac{1-\delta q}{2}=-n,n=0,1,2...,$ (31)
From which we obtain the energy as
$\mathcal{E}_{n}=-\frac{\hbar^{2}\alpha^{2}}{2\mu}\left(q\sqrt{\frac{2\mu\mathcal{D}_{e}}{\hbar^{2}\alpha^{2}}}-\left(n+\frac{1}{2}\right)\right)^{2},$ (32)
where $n=0,1,2,3,...,n_{\text{max}}$, and $n_{\text{max}}$ can be derived by $\frac{d\mathcal{E}_{n}}{dn}=0$, thus we obtain: $n_{\text{max}}\leq(\lambda q-1/2)$. For completeness, the wave function of the q-deformed Morse potential can be written as:
$\psi(y)=\mathcal{N}_{n}^{\delta}\text{e}^{-\frac{y}{2}}y^{s}L_{n}^{2s}(y),$ (33)
where $L_{n}^{2s}(y)$ are the associated Laguerre functions and $\mathcal{N}_{n}^{\delta}$ is the normalization constant.
Appendix B Quantum Otto Cycle: $\mathcal{Q}^{\text{QO}}_{\text{h}}$, $\mathcal{Q}^{\text{QO}}_{\text{c}}$ and $\mathcal{W}^{\text{QO}}$
Here, we present the expressions of the thermodynamic quantities used for the Otto engine performance analysis. Employing the energy equation and Eq. (15), we obtain the analytical expression for $\mathcal{Q}^{\text{QO}}_{\text{h}}$ as follows:
$\mathcal{Q}^{\text{QO}}_{\text{h}}=\frac{\gamma_{1}^{\text{h}}e^{\Lambda_{1}^{\text{h}}}}{\sqrt{\pi}\tau^{h}}+\frac{\Lambda_{3}^{\text{ch}}}{2\beta_{\text{h}}}-\frac{\xi_{\text{h}}^{2}}{2\beta_{\text{c}}\xi_{\text{c}}^{2}}+\frac{\Lambda_{2}^{\text{c}}e^{\Lambda_{1}^{\text{c}}}}{2\sqrt{\pi\beta_{\text{c}}\xi_{\text{c}}^{2}(-p)\tau^{c}}}$ (34)
where
$\tau^{c(h)}=\beta_{\text{c(h)}}\operatorname{erfc}\left(\gamma_{1}^{\text{c(h)}}\right),$ (35)
$\Lambda_{1}^{\text{c}}=\frac{1}{4}\beta_{\text{c}}\xi_{\text{c}}^{2}p(1-2\lambda_{\text{c}}q)^{2},$ (36)
$\Lambda_{1}^{\text{h}}=\frac{1}{4}\beta_{\text{h}}\xi_{\text{h}}^{2}p(1-2\lambda_{\text{h}}q)^{2},$ (37)
$\Lambda_{2}^{\text{ch}}=\xi_{\text{h}}^{2}p(2q(\lambda_{\text{c}}-2\lambda_{\text{h}})+1),$ (38)
$\Lambda_{3}^{\text{ch}}=2\beta_{\text{h}}\xi_{\text{h}}^{2}pq^{2}(\lambda_{\text{c}}-\lambda_{\text{h}})^{2}+1,$ (39)
$2\gamma_{1}^{\text{h}}=(1-2\lambda_{\text{h}}q)\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}$ (40)
Similarly, employing the energy equation and Eq. (16), we obtain the analytical expression for $\mathcal{Q}^{\text{QO}}_{\text{c}}$ as follows:
$\mathcal{Q}^{\text{QO}}_{\text{c}}$ $=$ $\frac{\gamma_{1}^{\text{c}}e^{\Lambda_{1}^{\text{c}}}}{\sqrt{\pi}\tau^{c}}+\frac{\Lambda_{2}^{\text{hc*}}e^{\Lambda_{1}^{\text{h}}}}{2\sqrt{\pi}\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}\text{erfc}\left(\gamma_{1}^{\text{h}}\right)}$ (42)
$+$ $\left(\frac{1}{2\beta_{\text{c}}}+\frac{\Lambda_{4}^{ch}}{2}\right)$
where
$\Lambda_{2}^{\text{hc*}}=\xi_{\text{c}}^{2}p(2q(\lambda_{\text{h}}-2\lambda_{\text{c}})+1),$ (43)
$2\gamma_{1}^{\text{c}}=(1-2\text{$\lambda$c}q)\sqrt{\beta_{\text{c}}\xi_{\text{c}}^{2}(-p)}$ (44)
$\gamma_{1}^{\text{h}}=\frac{1}{2}(1-2\lambda_{\text{h}}q)\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}$ (45)
$\Lambda_{4}^{ch}=\xi_{\text{c}}^{2}\left(2pq^{2}(\lambda_{\text{c}}-\lambda_{\text{h}})^{2}-\frac{1}{\beta_{\text{h}}\xi_{\text{h}}^{2}}\right)$ (46)
Similarly, using Eq. (17), we obtain the work output as follows:
$\mathcal{W}^{\text{QO}}=\frac{1}{2\sqrt{\pi}}\left(\frac{\eta+\Lambda_{11}}{\tau^{h}\beta_{\text{c}}\xi_{\text{h}}^{2}}-\frac{p\left(\Lambda_{12}+\Lambda_{13}\right)}{\left(\xi_{\text{c}}^{2}(-p)\right)^{3/2}\tau^{c}}\right)$ (47)
where
$\eta=\frac{\Lambda_{10}}{\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}},$ (48)
$\Lambda_{5}^{\text{ch}}=2\beta_{\text{c}}\xi_{\text{c}}^{2}pq^{2}(\lambda_{\text{c}}-\lambda_{\text{h}})^{2}-1,$ (49)
$\Lambda_{6}^{\text{ch}}=\xi_{\text{h}}^{2}(2\lambda_{\text{c}}q-4\lambda_{\text{h}}q+1)+\xi_{\text{c}}^{2}(2\lambda_{\text{c}}q-1)$ (50)
$\Lambda_{7}^{\text{ch}}=\xi_{\text{c}}^{2}(-4\lambda_{\text{c}}q+2\lambda_{\text{h}}q+1)+\xi_{\text{h}}^{2}(2\lambda_{\text{h}}q-1)$ (51)
$\Lambda_{8}^{\text{ch}}=\xi_{\text{h}}^{2}\left(\beta_{\text{c}}+\beta_{\text{h}}+2\beta_{\text{c}}\beta_{\text{h}}\xi_{\text{c}}^{2}pq^{2}(\lambda_{\text{c}}-\lambda_{\text{h}})^{2}\right)-\beta_{\text{c}}\xi_{\text{c}}^{2}$ (52)
$\Lambda_{9}^{\text{h}}=\frac{1}{2}\sqrt{\beta_{\text{h}}}\xi_{\text{h}}\sqrt{p}(1-2\lambda_{\text{h}}q)$ (53)
$\Lambda_{9}^{\text{c}}=\frac{1}{2}\sqrt{\beta_{\text{c}}}\xi_{\text{c}}\sqrt{p}(1-2\lambda_{\text{c}}q)$ (54)
$\Lambda_{10}=\sqrt{\pi}\sqrt{\beta_{\text{h}}}\xi_{\text{h}}\sqrt{p}\,\text{erfi}\left(\Lambda_{9}^{\text{h}}\right)\left(\Lambda_{8}^{\text{ch}}\right)$ (55)
$\Lambda_{11}=\sqrt{\pi}\left(\Lambda_{8}^{\text{ch}}\right)-\beta_{\text{c}}\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}e^{\Lambda_{1}^{\text{h}}}\left(\Lambda_{7}^{\text{ch}}\right)$ (56)
$\Lambda_{12}=\sqrt{\pi}\sqrt{\beta_{\text{c}}}\xi_{\text{c}}\xi_{\text{h}}^{2}\sqrt{p}\,\text{erfi}\left(\Lambda_{9}^{\text{c}}\right)\left(\Lambda_{5}^{\text{ch}}\right)$ (57)
$\Lambda_{13}=\sqrt{\pi}\xi_{\text{h}}^{2}\sqrt{\beta_{\text{c}}\xi_{\text{c}}^{2}(-p)}\left(\Lambda_{5}^{\text{ch}}\right)+\beta_{\text{c}}\xi_{\text{c}}^{2}p\,e^{\Lambda_{1}^{\text{c}}}\left(\Lambda_{6}^{\text{ch}}\right)$ (58)
Finally, using Eq. (18) and (47), we obtain an explicit analytical expression for the efficiency as follows:
$\eta^{{\text{QO}}}=\frac{\frac{\eta+\Lambda_{11}}{\beta_{\text{c}}\xi_{\text{h}}^{2}\tau^{h}}-\frac{p\left(\Lambda_{12}+\Lambda_{13}\right)}{(\beta_{\text{c}}\xi_{\text{c}}^{2}(-p))^{3/2}\text{erfc}\left(\gamma_{1}^{\text{c}}\right)}}{\frac{2\gamma_{1}^{\text{h}}e^{\Lambda_{1}^{\text{h}}}}{\tau^{h}}+\sqrt{\pi}\left(-\frac{\xi_{\text{h}}^{2}}{\beta_{\text{c}}\xi_{\text{c}}^{2}}+\frac{\Lambda_{3}^{\text{ch}}}{\beta_{\text{h}}}\right)+\frac{\Lambda_{2}^{\text{c}}e^{\Lambda_{1}^{\text{c}}}}{\sqrt{\beta_{\text{c}}\xi_{\text{c}}^{2}(-p)}\text{erfc}\left(\gamma_{1}^{\text{c}}\right)}}$ (59)